OpenStudy (anonymous):

how to simplify: (x^3\4)^5

OpenStudy (anonymous):

take (x^3)^5 which becomes X^8 and then do 4^5 and you get (x^8)/(4^8)

OpenStudy (unklerhaukus):

is that \[(x^3/4)^5\],\[(x^{3/4})^5\]or something else?

OpenStudy (anonymous):

no youre making it harder just take the numerator to the fifth power and then the denominator to the fifth

OpenStudy (unklerhaukus):

\[(a^n)^m=a^{n\times m}\]

OpenStudy (anonymous):

\[\frac{ x^8 }{ 4^5 }\]

OpenStudy (unklerhaukus):

that is not right

OpenStudy (anonymous):

unless you are saying x is raised to the (3/4)

OpenStudy (unklerhaukus):

\[(x^3/4)^5=(xxx/4)^5\\=(xxx/4)(xxx/4)(xxx/4)(xxx/4)(xxx/4)\\(xxxxxxxxxxxxxxx/4\times4\times4\times4\times4)\]

OpenStudy (anonymous):

okay I see now yes that is right

OpenStudy (unklerhaukus):

the question as stated by OP is unclear \ (backslash) isn't usually used for fractions